A boat takes 4 hours to travel 37.09 km down a river, then 6.6 hours to return. How fast is the river flowing?
Answer #1

in first condition boat and the river flow are in the same direction so the relative velocity is

V = Vb + Vr

and hence

relative velocity = distance / time

( Vb + Vr ) = 37.09/4

Vb + Vr = 9.27............................(1)

in the second condition velocities are in opposite direction

hence  V = Vb  - Vr

Vb  - Vr = 37.09 / 6.6

Vb  - Vr = 5.62 ..............................(2)

by subtracting equation 2 from 1 we get

2Vr = 3.65

so Vr = 1.825

Hence the river flowing with 1.825 km / h